3. Calculate the pH of a solution that contains 0.250 M HCOOH (formic acid) and 0.100 M NaCOOH (sodium formate). Given Ka = 1.8 x 10 - 4 for HCOOH. Please include the ICE table.
Answer #1

HCOOH Ka = 1.8 x 10 - 4

pKa = -log Ka = -log (1.8 x 10 - 4)

       = 3.74

HCOOH + NaOH ------------------> NaCOOH + H2O

0.35             0.1                                    0               0     --------------- initial

-0.1             -0.1                                   +0.1          +0.1 ------------ change

0.25              0                                      0.1             0.1     ----------- equilibrium

pH = pKa + log [salt / acid]

      = 3.74 + log [0.1 / 0.250]

      = 3.34

pH = 3.34